Question 7 to 10 solve it in your answer sheet and scan it and upload as pdf. Students cannot type or send answer separately. Students can scan and upload only as pdf or word Document


7) In balance diet pie-chart what is the Total angle drawn for Fruits and vegetables, Protein and

Fibre-rich carbohydrates?

8) In above survey How many people eat fast food daily?

9 ) If 900 people were surveyed, How many people never eat fast food in out – lets?

10) If 900 people were surveyed, How many people eat fast food two times in a week?


1
Expert's answer
2021-02-24T10:24:25-0500

"\displaystyle\textsf{Angle for fruits and vegetables} = \frac{40}{100} \times 360\degree = 144\degree\\


\textsf{Angle for protein} = \frac{25}{100} \times 360\degree = 90\degree\\

\textsf{Angle for fibre-rich carbohydrates} = \frac{25}{100} \times 360\degree = 90\degree"